).Given a 0.1M aqueous solution each of NH4Cl and CH3COONa (ammonium ethanoate),state and explain which of them you would expect to have a lower pH


1
Expert's answer
2020-07-07T14:36:17-0400

NH4Cl is the salt of strong acid (HCl) and weak base (NH4OH)

pH of NH4Cl is = 7-1/2pKb – 1/2logC      [ pKb of NH4OH= 4.75, C= 0.1 M]

                        = 7- (4.75/2) -1/2log0.1

                      = 5.162

CH3COONa is the salt of weak acid (CH3COOH) and strong base (NaOH)

pH of CH3COONa  is = 7+ 1/2pKa + 1/2logC      [ pKa of CH3COOH = 4.76, C= 0.1 M]

                        = 7+(4.76/2) +1/2log0.1

                      = 8.88

ammonium ethanoate is the salt of weak acid (CH3COOH) and weak base (NH4OH)

pH of ammonium ethanoate is = 7+1/2pKa – 1/2pKb

                                               = 7+(4.76/2) - (4.75/2)

                                                 = 7.005

NH4Cl is the lower pH (Answer)